The formula C=59(F−32) expresses the relationship between Fahrenheit​ temperature, F, and Celsius​ temperature, C. Use the formula to convert 59°F to its equivalent temperature on the Celsius scale.



1
Expert's answer
2022-02-01T17:16:38-0500

"C= \frac {5}{9}(F-32)"

When "F=59"


"C=\frac{5}{9}(59-32)=15°c"